2.When a consumer’s disposable income increases from 600 TL to 800 TL , her level of personal consumption expenditures increases from 350 TL to 400 TL. What is her marginal propensity to save (MPS)?

a.

0.50

b.

0.25

c.

0.75

d.

0.33.

3.Which of the following correctly fills in the three blanks in the description given below?

The speculative demand for money schedule flattens out at a very ……......…... rate of interest, showing that at these rates, there is a general expectation of capital losses on bonds that outweigh interest earnings. At this rate, increments to wealth would be held in the form of ………......……., with no further drop in the interest rate. Keynes termed this situation the .................

a.

high ; bond ; capital dilution

b.

high ; money ; capital dilution

c.

low ; bond ; liquidity trap

d.

low ; money ; liquidity trap


1
Expert's answer
2021-07-07T13:41:13-0400

2."MPC =\frac {dS} {dY}"


"MPC=\frac {400-350}{800-600}"


MPC=0.25


"1-MPC =MPS"

"1-0.25=0.75"

MPS =0.75


3.high ; bond ; capital dilution